OpenStudy (msmr):

Jeremy can paint a house in 3 days, Katie in 5 days and Bill in 7 days. How long will it take them to paint the house?

OpenStudy (saifoo.khan):

Jermy's share = 1/3 Katie's share = 1/5 Bill's share = 1/7 therefore, \[\frac 1x = \frac 13 + \frac 15 + \frac 17\]

OpenStudy (msmr):

thank you! i wasn't sure whether it would be the fraction or not.

OpenStudy (saifoo.khan):

Welcome.